OpenStudy (twistrose):

During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form.

OpenStudy (twistrose):

i dont get it

OpenStudy (antiprime):

its 0.318

OpenStudy (antiprime):

it wont let me delete the long text

OpenStudy (antiprime):

i fixed it

OpenStudy (antiprime):

so what i would so is but 12_3/4 as a deci al +12.75 then divide that by 40%

OpenStudy (twistrose):

which is 0.40

OpenStudy (texaschic101):

u dont divide

OpenStudy (texaschic101):

12.75 - 0.4 (12.75) = ?

OpenStudy (twistrose):

157.4625

OpenStudy (texaschic101):

12.75 - .4 (12.75) = 12.75 - 5.1 = 7.65 = 7 13/20

OpenStudy (twistrose):

ok thanks